OpenStudy (angelax0):

The function g is defined by g(x)=2x^2-4 Find g(5x).

OpenStudy (welshfella):

replace the x in the expression for g(x) by 5x and simplify.

OpenStudy (angelax0):

2(5x)^2-4 = 100-4?

OpenStudy (welshfella):

the left side of your answer is correct but where as the 'x' gone?

OpenStudy (angelax0):

5x = 2(5x)^2 - 4

OpenStudy (welshfella):

g(5x) = 2(5x)^2 - 4

OpenStudy (angelax0):

how would you solve this?

OpenStudy (welshfella):

now simplify

OpenStudy (angelax0):

g(5x)=100-4 = g(5x)=96 ?

OpenStudy (welshfella):

No g)5x) = 2* 25x^2 - 4 = 50x^2 - 4 we are not solving an equation for x Just simplifying 2(5x)^2 - 4

OpenStudy (angelax0):

So that would be the final answer? g(5x)=2(5x)^2-4 ?

OpenStudy (welshfella):

no see my last post 50x^2 - 4 is the final answer

OpenStudy (angelax0):

So you wouldnt do the 50x^2? you would just leave it like that?

OpenStudy (mathmale):

You must keep the function name, g(x). Write out g(x). Wherever you see 'x,' write '5x' instead.

OpenStudy (mathmale):

Write out g(x). Next, replace every 'x' in this function with '5x.' Keep the label 'g.' The new function is then g(5x). Write it out.